C=300+0.75Y, I=400, G=300,T=100

determine the equilibrium level of the income using expenditure and injection-leakage approach


1
Expert's answer
2022-03-09T11:02:45-0500

Solution

"C=300+0.75Y"

"I=400"

"G=300"

"T=100"


The expenditure approach method is given by;

"Y=C+I+G+(X-M)"


Therefore;

"Y=300+0.75(Y-100)+400+300"

"Y=300+0.75Y-75+400+300"

Put like terms together and simplify

"Y-0.75Y=300-75+400+300"

"0.25Y=925"

"Y=\frac{925}{0.25}"

"Y=3,700"